OpenStudy (anonymous):

Suppose f(x)=-3x+2 and g(x)=1/2x-1. Find each value. 1. f(1/3) 2. 3g(4)

OpenStudy (anonymous):

plug in 1/3 for f(x), so f(1/3)=-3(1/3)+2, which equals 1 3g(4)= 1/2(4)-1, =13(1)=3

OpenStudy (anonymous):

* it is not 13, the one and three are separate 3(1)=3 its not 13(1)